# On the Fragility of Active Learners

### v2

@@ 0.A.1 Datasets
Dataset
# classes
Description
sst-2
2
Single sentences extracted from movie reviews with their sentiment label [30].
imdb
2
Movie reviews with corresponding sentiment label [19].
agnews
4
News articles with their topic category [36].
pubmed
4
Sentences in medical articles’ abstracts which are labeled with their role on the abstract [8].
dbpedia-5
5
A subset of dbpedia [36] which contains Wikipedia articles accompanied by a topic label. The original dataset’s instances are evenly distributed across $14$ classes. To form dbpedia-5, we use only the first $5$ classes: Company, EducationalInstitution, Artist, Athlete, OfficeHolder. This was done to reduce the training time of one-vs-all classifiers, e.g., LinearSVC.
@@ 0.A.3 Reproducibility Experiments
AL
Dataset
AL loop
Classifier & text representation
QS parameters
Metric
CAL
agnews
b=2280
@@
# neighbors=10
F1-macro
DAL
cola
b=50
@@ 0.A.4.2 Hyperparameters search for prediction pipelines
Classifier
Hyperparameters
RoBERTa
roberta-base

### v3

@@ 0.A.1 Datasets
Dataset # classes Description
sst-2 2 Single sentences extracted from movie reviews with their sentiment label [30].
imdb 2 Movie reviews with corresponding sentiment label [19].
agnews 4 News articles with their topic category [36].
pubmed 4 Sentences in medical articles’ abstracts which are labeled with their role on the abstract [8].
dbpedia-5 5 A subset of dbpedia [36] which contains Wikipedia articles accompanied by a topic label. The original dataset’s instances are evenly distributed across $14$ classes. To form dbpedia-5, we use only the first $5$ classes: Company, EducationalInstitution, Artist, Athlete, OfficeHolder. This was done to reduce the training time of one-vs-all classifiers, e.g., LinearSVC.
@@ 0.A.3 Reproducibility Experiments
AL Dataset AL loop Classifier & text representation QS parameters Metric
CAL agnews b=2280
@@
# neighbors=10 F1-macro
DAL cola b=50
@@ 0.A.4.2 Hyperparameters search for prediction pipelines
Classifier Hyperparameters
RoBERTa roberta-base